AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 01:28:11 2025
Author
*
:
Subject
*
:
<p>22.00</p> <p>Carsten Brandau stopping e-target experiment Unattended beam overnight. No operators watching.</p> <p>10^5 (nominal) 20Ne7+ at 380keV/u beam available</p> <p>Carsten reports beam is unstable, often disappearing, and in bunched mode.</p> <p>4x MSL type TTT14 DSSSDs not moving, locked in out position</p> <p>Starting CARME MIDAS DAQ ASIC settings 2024Apr24-14.14.03</p> <p>slow comparator 0x64 E</p> <p>G&G Ortec 448 Pulser OFF</p> <p>DSSSD bias & leakage currents - attachment 1</p> <p>FEE64 temperatures OK - attachment 2</p> <p>All system wide checks OK, WR timestamps OK - attachment 3</p> <p>ADC data item stats - attachment 4</p> <p>Rates dominated by two hot channels aida15 per FEE64 Stat spectra - attachment 5</p> <p>per FEE64 1.8.L spectra - attachments 6-7 no obvious beam energy peak, continuum of energies from below beam energy to threshold</p> <p>per FEE64 1.8.W spectra - attachments 8-9</p> <p> </p> <p>22.55 Start data file APR24/R5 Merger, TapeServer etc - attachments 10-12</p> <p>270k data items/s</p> <p> </p> <p>23:29 Activated CARMESort script. See attachments 13-15.</p> <p>No obvious beam or signal of any kind. Temperatures OK.</p> <p> </p> <p>00:13 DAQ continues OK. System-wide check OK. Temps OK. Still no clear sign of beam.</p> <p> </p> <p>01:03 Daq continues OK. System wide checks OK, Temps OK.</p> <p>Only low energies observed in the energy spectrum. Looks a bit like the low energy noise observed in our previous runs with no/very few counts above what would be beam energy (~7.6 MeV)</p> <p>Note the xy plots are not all orientated correctly (strips are correct in the sort code but root axis need to be flipped - I can;t remember how atm). The part of the plots with counts are those closest to the beam. </p> <p>02:00 DAQ continues OK. System wide checks OK. Temps OK.</p> <p>Spectra still looks the same. I have noticed the right hand side rate is higher than the left. EyEx cut spectrum also looks like much more counts outside of the cut.</p> <p>03:40</p> <p>DAQ continues. All checks and temps OK.</p> <p>I can now see a small peak at 5 MeV in the bottom left dssd. This dssd also has the least number of counts in the low energy continuum.</p> <p>I have logged back into grafana on x2go to get the number ions in the ring. Attachment 17</p> <p>04:23 (05:23 EU)</p> <p>DAQ stopped, merger stopped, tape server stopped. Bias OFF. FEEs turned OFF. Screen shots taken of the monitoring code plots - not yet uploaded.</p> <p>Windows and servers remain up in case we want to do some more tests in the near future.</p> <p>Attachment 18 shows beam intensity for total run time from Grafana</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown